# Mean-Field Sampling for Cooperative Multi-Agent Reinforcement Learning

### v1

@@ Appendix D Proof of Lipschitz-continuity Bound
since $\sum_{(s_{g}^{\prime},s_{\Delta}^{\prime})\in\mathcal{S}_{g}\times\mathcal{S}_{l}^{|\Delta|}}\mathcal{J}_{k}[s_{g}^{\prime},s_{\Delta}^{\prime},s_{g},a_{g},s_{\Delta},a_{\Delta}]\mathcal{J}_{k}[s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},s_{g}^{\prime},a_{g},s_{g}^{\prime},a_{\Delta}]$ is the stochastic probability function corresponding to the two-step evolution of the joint states from $(s_{g},s_{\Delta})$ to $(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})$ under actions $a_{g},a_{\Delta}$. This can be thought of as an analogously to the fact that a $1$ on the $(i,j)$’th entry on the square of a $0/1$ adjacency matrix of a graph represents the fact that there is a path of distance $2$ between vertices $i$ and $j$. Finally, the third equality recovers the definition of the expectation, with respect to the joint probability function $\mathcal{J}_{k}^{2}$. ∎
We next show that $\lim_{T\to\infty}\mathbb{E}_{\Delta\in{[n]\choose k}}\widehat{Q}_{k}^{T}(s_{g},s_{\Delta},a_{g},a_{\Delta})=\mathbb{E}_{\Delta\in{[n]\choose k}}\widehat{Q}_{k}^{*}(s_{g},s_{\Delta},a_{g},a_{\Delta})=Q^{*}(s,a)$.
Lemma D.11.
$Q^{*}(s,a)-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\widehat{Q}_{k}^{T}(s_{g},s_{\Delta},a_{g},a_{\Delta})\leq\gamma^{T}\cdot\frac{\widetilde{r}}{1-\gamma}$
@@
We bound the differences between $\widehat{Q}_{k}^{T}$ at each Bellman iteration of our approximation to $Q^{*}$.
@@
$Q^{*}(s,a)-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\widehat{Q}_{k}^{T}(s_{g},s_{\Delta},a_{g},a_{\Delta})$
$=\mathcal{T}Q^{*}(s,a)-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\widehat{\mathcal{T}}_{k}\widehat{Q}_{k}^{T-1}(s_{g},s_{\Delta},a_{g},a_{\Delta})$
@@
$\quad\quad\quad\quad-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\bigg{[}r_{\Delta}(s_{g},s_{\Delta},a_{g},a_{\Delta})+\gamma\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim{P}_{g}(\cdot|s_{g},a_{g})\\ s_{i}^{\prime}\sim{P}_{l}(\cdot|s_{i},a_{i}s_{g}),\forall i\in\Delta\end{subarray}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}}Q_{k}^{T}(s_{g}^{\prime},s_{\Delta}^{\prime},a_{g}^{\prime},a_{\Delta}^{\prime})\bigg{]}$
Next, observe that $r_{[n]}(s_{g},s_{[n]},a_{g},a_{[n]})=\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}r_{[\Delta]}(s_{g},s_{\Delta},a_{g},a_{\Delta})$. To prove this, we write:
@@
$\sup_{(s,a)\in\mathcal{S}\times\mathcal{A}}\bigg{[}Q^{*}(s,a)-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\widehat{Q}_{k}^{T}(s_{g},s_{[n]},a_{g},a_{\Delta})\bigg{]}$
$=\sup_{(s,a)\in\mathcal{S}\times\mathcal{A}}\bigg{[}\mathcal{T}Q^{*}(s,a)-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\widehat{\mathcal{T}}_{k}\widehat{Q}_{k}^{T-1}(s_{g},s_{\Delta},a_{g},a_{\Delta})\bigg{]}$
$=\gamma\sup_{(s,a)\in\mathcal{S}\times\mathcal{A}}\bigg{[}\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim{P}(\cdot|s_{g},a_{g}),\\ s_{i}^{\prime}\sim P_{l}(\cdot|s_{i},a_{i},s_{g}),\forall i\in[n]\end{subarray}}\max_{a^{\prime}\in\mathcal{A}}Q^{*}(s^{\prime},a^{\prime})-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim{P}_{g}(\cdot|s_{g},a_{g}),\\ s_{i}^{\prime}\sim{P}_{l}(\cdot|s_{i},a_{i},s_{g}),\forall i\in\Delta\end{subarray}}\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},\\ a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}\end{subarray}}\widehat{Q}_{k}^{T-1}(s_{g}^{\prime},s_{\Delta}^{\prime},a_{g}^{\prime},a_{\Delta}^{\prime})\bigg{]}$
$=\gamma\sup_{(s,a)\in\mathcal{S}\times\mathcal{A}}\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim{P}_{g}(\cdot|s_{g},a_{g}),\\ s_{i}^{\prime}\sim P_{l}(\cdot|s_{i},a_{i},s_{g}),\forall i\in[n]\end{subarray}}\bigg{[}\max_{a^{\prime}\in\mathcal{A}}Q^{*}(s^{\prime},a^{\prime})-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}\end{subarray}}\widehat{Q}_{k}^{T-1}(s_{g}^{\prime},s_{\Delta}^{\prime},a_{g}^{\prime},a_{\Delta}^{\prime})\bigg{]}$
$\leq\gamma\sup_{(s,a)\in\mathcal{S}\times\mathcal{A}}\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim{P}_{g}(\cdot|s_{g},a_{g}),\\ s_{i}^{\prime}\sim P_{l}(\cdot|s_{i},a_{i},s_{g}),\forall i\in[n]\end{subarray}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{[n]}^{\prime}\in\mathcal{A}_{l}^{n}}\bigg{[}Q^{*}(s^{\prime},a^{\prime})-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\widehat{Q}_{k}^{T-1}(s_{g}^{\prime},s_{\Delta}^{\prime},a_{g}^{\prime},a_{\Delta}^{\prime})\bigg{]}$
$\leq\gamma\sup_{(s^{\prime},a^{\prime})\in\mathcal{S}\times\mathcal{A}}\bigg{[}Q^{*}(s^{\prime},a^{\prime})-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\widehat{Q}_{k}^{T-1}(s^{\prime}_{g},s^{\prime}_{\Delta},a^{\prime}_{g},a^{\prime}_{\Delta})\bigg{]}$
@@
$Q^{*}(s,a)-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\widehat{Q}_{k}^{T}(s_{g},s_{\Delta},a_{g},a_{\Delta})$ $\leq\gamma^{T}\sup_{(s^{\prime},a^{\prime})\in\mathcal{S}\times\mathcal{A}}\bigg{[}Q^{*}(s^{\prime},a^{\prime})-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\widehat{Q}_{k}^{0}(s^{\prime}_{g},s^{\prime}_{\Delta},a^{\prime}_{g},a^{\prime}_{\Delta})\bigg{]}$
@@
$=\gamma^{T}\cdot\frac{\widetilde{r}}{1-\gamma}$
The first inequality follows from the $\gamma$-contraction property of the update procedure, and the ensuing equality follows from our bound on the maximum possible value of $Q$ from Lemma B.5 and noting that $\widehat{Q}_{k}^{0}:=0$. Therefore, as $T\to\infty$,
$Q^{*}(s,a_{g})-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\widehat{Q}_{k}^{T}(s_{g},s_{\Delta},a_{g})\to 0,$
@@
Corollary D.11.1.
Since $\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\widehat{Q}_{k}^{T}(s_{g},s_{\Delta},a_{g},a_{\Delta}):=\mathbb{E}_{\Delta\in{[n]\choose k}}\widehat{Q}_{k}^{T}(s_{g},s_{\Delta},a_{g},a_{\Delta})$, we therefore get:
$Q^{*}(s,a)-\lim_{T\to\infty}\mathbb{E}_{\Delta\in{[n]\choose k}}\widehat{Q}_{k}^{T}(s_{g},s_{\Delta},a_{g},a_{\Delta})\leq\lim_{T\to\infty}\gamma^{T}\cdot\frac{\widetilde{r}}{1-\gamma}$
@@
$Q^{*}(s,a)-\mathbb{E}_{\Delta\in{[n]\choose k}}\widehat{Q}_{k}^{*}(s_{g},s_{\Delta},a_{g},a_{\Delta})\leq 0$
@@
$Q^{*}(s,a)-\mathbb{E}_{\Delta\in{[n]\choose k}}\widehat{Q}_{k}^{*}(s_{g},s_{\Delta},a_{g},a_{\Delta})=0,$
since $\widehat{Q}_{k}^{*}(s_{g},s_{\Delta},a_{g},a_{\Delta})\leq Q^{*}(s,a)$.
Lemma D.12.
The absolute difference between the expected maximums between $\widehat{Q}_{k}$ and $\widehat{Q}_{k^{\prime}}$ is atmost the maximum of the absolute difference between $\widehat{Q}_{k}$ and $\widehat{Q}_{k^{\prime}}$, where the expectations are taken over any joint distributions of states $\mathcal{J}$, and the maximums are taken over the actions.
$\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot,\cdot|s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}\bigg{[}\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},\\ a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}\end{subarray}}\widehat{Q}_{k}^{T}(s_{g}^{\prime},s_{\Delta}^{\prime},a_{g}^{\prime},a_{\Delta}^{\prime})$ $-\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},\\ a_{\Delta^{\prime}}^{\prime}\in\mathcal{A}_{l}^{k^{\prime}}\end{subarray}}\widehat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime},a_{g}^{\prime},a^{\prime}_{\Delta^{\prime}})\bigg{]}\bigg{|}$
$\leq\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},\\ a^{\prime}_{\Delta\cup\Delta^{\prime}}\in\mathcal{A}_{l}^{|\Delta\cup\Delta^{\prime}|}\end{subarray}}\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot,\cdot|s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}\bigg{[}\widehat{Q}_{k}^{T}(s_{g}^{\prime},s_{\Delta}^{\prime},a_{g}^{\prime},a_{\Delta}^{\prime})$ $-\widehat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime},a_{g}^{\prime},a_{\Delta^{\prime}}^{\prime})\bigg{]}\bigg{|}$
@@
$a_{g}^{*},a_{\Delta}^{*}:=\arg\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},\\ a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}\end{subarray}}\widehat{Q}_{k}^{T}(s_{g}^{\prime},F_{s^{\prime}_{\Delta}},a_{g}^{\prime},a_{\Delta}^{\prime})$
$\widetilde{a}_{g}^{*},\widetilde{a}_{\Delta^{\prime}}^{*}:=\arg\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},\\ a_{\Delta^{\prime}}^{\prime}\in\mathcal{A}_{l}^{k^{\prime}}\end{subarray}}\widehat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},F_{s^{\prime}_{\Delta^{\prime}}},a_{g}^{\prime},a_{\Delta^{\prime}}^{\prime})$

### v3

@@ Appendix E Proof of Lipschitz-continuity Bound
since $\sum_{(s_{g}^{\prime},s_{\Delta}^{\prime})\in\mathcal{S}_{g}\times\mathcal{S}_{l}^{|\Delta|}}\mathcal{J}_{k}[s_{g}^{\prime},s_{\Delta}^{\prime},s_{g},a_{g},s_{\Delta},a_{\Delta}]\mathcal{J}_{k}[s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},s_{g}^{\prime},a_{g},s_{g}^{\prime},a_{\Delta}]$ is the stochastic probability function corresponding to the two-step evolution of the joint states from $(s_{g},s_{\Delta})$ to $(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})$ under actions $a_{g},a_{\Delta}$. This can be thought of as an analogously to the fact that a $1$ on the $(i,j)$’th entry on the square of a $0/1$ adjacency matrix of a graph represents the fact that there is a path of distance $2$ between vertices $i$ and $j$.
Finally, the third equality recovers the definition of the expectation, with respect to the joint probability function $\mathcal{J}_{k}^{2}$. ∎
We next show $\lim_{T\to\infty}\mathbb{E}_{\Delta\in{[n]\choose k}}\hat{Q}_{k}^{T}(s_{g},s_{\Delta},a_{g},a_{\Delta})=\mathbb{E}_{\Delta\in{[n]\choose k}}\hat{Q}_{k}^{*}(s_{g},s_{\Delta},a_{g},a_{\Delta})=Q^{*}(s,a)$.
Lemma E.12.
The $Q^{*}$ function is the average value of each of the (${n\choose k}$ sub-sampled $\hat{Q}^{*}_{k}$-functions.
$Q^{*}(s,a)-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\hat{Q}_{k}^{T}(s_{g},s_{\Delta},a_{g},a_{\Delta})\leq\gamma^{T}\cdot\frac{\tilde{r}}{1-\gamma}$
@@
We bound the differences between $\hat{Q}_{k}^{T}$ at each Bellman iteration of our approximation to $Q^{*}$.
@@
$Q^{*}(s,a)-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\hat{Q}_{k}^{T}(s_{g},s_{\Delta},a_{g},a_{\Delta})$
$=\mathcal{T}Q^{*}(s,a)-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\hat{\mathcal{T}}_{k}\hat{Q}_{k}^{T-1}(s_{g},s_{\Delta},a_{g},a_{\Delta})$
@@
$-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\bigg{[}r_{\Delta}(s_{g},s_{\Delta},a_{g},a_{\Delta})+\gamma\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim{P}_{g}(\cdot|s_{g},a_{g})\\ s_{i}^{\prime}\sim{P}_{l}(\cdot|s_{i},a_{i}s_{g}),\forall i\in\Delta\end{subarray}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}}Q_{k}^{T}(s_{g}^{\prime},s_{\Delta}^{\prime},a_{g}^{\prime},a_{\Delta}^{\prime})\bigg{]}$
Next, observe that $r_{[n]}(s_{g},s_{[n]},a_{g},a_{[n]})=\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}r_{[\Delta]}(s_{g},s_{\Delta},a_{g},a_{\Delta})$.
To prove this, we write:
@@
$\sup_{(s,a)\in\mathcal{S}\times\mathcal{A}}\bigg{[}Q^{*}(s,a)-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\hat{Q}_{k}^{T}(s_{g},s_{[n]},a_{g},a_{\Delta})\bigg{]}$
$=\sup_{(s,a)\in\mathcal{S}\times\mathcal{A}}\bigg{[}\mathcal{T}Q^{*}(s,a)-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\hat{\mathcal{T}}_{k}\hat{Q}_{k}^{T-1}(s_{g},s_{\Delta},a_{g},a_{\Delta})\bigg{]}$
$=\gamma\sup_{(s,a)\in\mathcal{S}\times\mathcal{A}}\bigg{[}\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim{P}(\cdot|s_{g},a_{g}),\\ s_{i}^{\prime}\sim P_{l}(\cdot|s_{i},a_{i},s_{g}),\forall i\in[n]\end{subarray}}\max_{a^{\prime}\in\mathcal{A}}Q^{*}(s^{\prime},a^{\prime})$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim{P}_{g}(\cdot|s_{g},a_{g}),\\ s_{i}^{\prime}\sim{P}_{l}(\cdot|s_{i},a_{i},s_{g}),\forall i\in\Delta\end{subarray}}\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},\\ a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}\end{subarray}}\hat{Q}_{k}^{T-1}(s_{g}^{\prime},s_{\Delta}^{\prime},a_{g}^{\prime},a_{\Delta}^{\prime})\bigg{]}$
$=\gamma\sup_{(s,a)\in\mathcal{S}\times\mathcal{A}}\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim{P}_{g}(\cdot|s_{g},a_{g}),\\ s_{i}^{\prime}\sim P_{l}(\cdot|s_{i},a_{i},s_{g}),\forall i\in[n]\end{subarray}}\bigg{[}\max_{a^{\prime}\in\mathcal{A}}Q^{*}(s^{\prime},a^{\prime})$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}\end{subarray}}\hat{Q}_{k}^{T-1}(s_{g}^{\prime},s_{\Delta}^{\prime},a_{g}^{\prime},a_{\Delta}^{\prime})\bigg{]}$
$\leq\gamma\sup_{(s,a)\in\mathcal{S}\times\mathcal{A}}\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim{P}_{g}(\cdot|s_{g},a_{g}),\\ s_{i}^{\prime}\sim P_{l}(\cdot|s_{i},a_{i},s_{g}),\forall i\in[n]\end{subarray}}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{[n]}^{\prime}\in\mathcal{A}_{l}^{n}}\bigg{[}Q^{*}(s^{\prime},a^{\prime})-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\hat{Q}_{k}^{T-1}(s_{g}^{\prime},s_{\Delta}^{\prime},a_{g}^{\prime},a_{\Delta}^{\prime})\bigg{]}$
$\leq\gamma\sup_{(s^{\prime},a^{\prime})\in\mathcal{S}\times\mathcal{A}}\bigg{[}Q^{*}(s^{\prime},a^{\prime})-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\hat{Q}_{k}^{T-1}(s^{\prime}_{g},s^{\prime}_{\Delta},a^{\prime}_{g},a^{\prime}_{\Delta})\bigg{]}$
@@
$Q^{*}(s,a)$ $-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\hat{Q}_{k}^{T}(s_{g},s_{\Delta},a_{g},a_{\Delta})$
$\leq\gamma^{T}\sup_{(s^{\prime},a^{\prime})\in\mathcal{S}\times\mathcal{A}}\bigg{[}Q^{*}(s^{\prime},a^{\prime})-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\hat{Q}_{k}^{0}(s^{\prime}_{g},s^{\prime}_{\Delta},a^{\prime}_{g},a^{\prime}_{\Delta})\bigg{]}$
@@
$=\gamma^{T}\cdot\frac{\tilde{r}}{1-\gamma}$
The first inequality follows from the $\gamma$-contraction property of the update procedure, and the ensuing equality follows from our bound on the maximum possible value of $Q$ from Lemma C.4 and noting that $\hat{Q}_{k}^{0}:=0$.
Therefore, as $T\to\infty$,
$Q^{*}(s,a_{g})-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\hat{Q}_{k}^{T}(s_{g},s_{\Delta},a_{g})\to 0,$
@@
Corollary E.13.
Since $\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\hat{Q}_{k}^{T}(s_{g},s_{\Delta},a_{g},a_{\Delta}):=\mathbb{E}_{\Delta\in{[n]\choose k}}\hat{Q}_{k}^{T}(s_{g},s_{\Delta},a_{g},a_{\Delta})$, we therefore get:
$Q^{*}(s,a)-\lim_{T\to\infty}\mathbb{E}_{\Delta\in{[n]\choose k}}\hat{Q}_{k}^{T}(s_{g},s_{\Delta},a_{g},a_{\Delta})\leq\lim_{T\to\infty}\gamma^{T}\cdot\frac{\tilde{r}}{1-\gamma}$
@@
$Q^{*}(s,a)-\mathbb{E}_{\Delta\in{[n]\choose k}}\hat{Q}_{k}^{*}(s_{g},s_{\Delta},a_{g},a_{\Delta})\leq 0$
@@
$Q^{*}(s,a)-\mathbb{E}_{\Delta\in{[n]\choose k}}\hat{Q}_{k}^{*}(s_{g},s_{\Delta},a_{g},a_{\Delta})=0,$
since $\hat{Q}_{k}^{*}(s_{g},s_{\Delta},a_{g},a_{\Delta})\leq Q^{*}(s,a)$.
Lemma E.14.
The absolute difference between the expected maximums between $\hat{Q}_{k}$ and $\hat{Q}_{k^{\prime}}$ is atmost the maximum of the absolute difference between $\hat{Q}_{k}$ and $\hat{Q}_{k^{\prime}}$, where the expectations are taken over any joint distributions of states $\mathcal{J}$, and the maximums are taken over the actions.
$\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot,\cdot|s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}\bigg{[}\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}\end{subarray}}\hat{Q}_{k}^{T}(s_{g}^{\prime},s_{\Delta}^{\prime},a_{g}^{\prime},a_{\Delta}^{\prime})$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta^{\prime}}^{\prime}\in\mathcal{A}_{l}^{k^{\prime}}\end{subarray}}\hat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime},a_{g}^{\prime},a^{\prime}_{\Delta^{\prime}})\bigg{]}\bigg{|}$
$\leq$ $\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},a^{\prime}_{\Delta\cup\Delta^{\prime}}\in\mathcal{A}_{l}^{|\Delta\cup\Delta^{\prime}|}\end{subarray}}\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot,\cdot|s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}\bigg{[}\hat{Q}_{k}^{T}(s_{g}^{\prime},s_{\Delta}^{\prime},a_{g}^{\prime},a_{\Delta}^{\prime})$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\hat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime},a_{g}^{\prime},a_{\Delta^{\prime}}^{\prime})\bigg{]}\bigg{|}$
@@
$a_{g}^{*},a_{\Delta}^{*}:=\arg\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},\\ a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}\end{subarray}}\hat{Q}_{k}^{T}(s_{g}^{\prime},F_{s^{\prime}_{\Delta}},a_{g}^{\prime},a_{\Delta}^{\prime})$
$\tilde{a}_{g}^{*},\tilde{a}_{\Delta^{\prime}}^{*}:=\arg\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},\\ a_{\Delta^{\prime}}^{\prime}\in\mathcal{A}_{l}^{k^{\prime}}\end{subarray}}\hat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},F_{s^{\prime}_{\Delta^{\prime}}},a_{g}^{\prime},a_{\Delta^{\prime}}^{\prime})$